Let demand be given by Q = 150 - P + 2Y. This is the same...

  1. Let demand be given by Q = 150 - P + 2Y. This is the same for all problems of this type. Let r = 10%. Let Y = 50 across both periods. Let MC = 0. Let reserves = 200. Consider the basic two-period model. What is consumption of the resource in the present IF THE OWNER OF THE RESOURCE IS A MONOPOLIST?

    94.29

    98.81

    101.19

    105.71

    none of the above
Answer #1

Answer option 3) 101.19

Q= 150-P+2*50

Q= 250-P

P= 250-Q

MR = 250-2*Q

so at eqm extraction rule of monopoly

250-2Q1 = (250-Q2)/(1+r)

& From resource constraint : Q1+Q2= 200

Q2= 200-Q1

so,250-2Q1 = (250-200+Q1)/1.1

solving Q1*= 101.19